OpenStudy (anonymous):

Ally is wearing a witch’s costume for Halloween. Her hat is in the shape of a cone with a diameter of 9 inches and a height of 14 inches. What is the volume of Ally’s hat? Use for pi. 382 in3 297 in3 1,188 in3 891 in3

OpenStudy (anonymous):

do you know the formula of a cone

OpenStudy (anonymous):

do you know how cone looks like? :)

OpenStudy (anonymous):



OpenStudy (anonymous):

its volume formula is: \[\LARGE V=\frac13 \pi \cdot r^2\cdot h\]

OpenStudy (anonymous):

okayyy now whaaat?

OpenStudy (anonymous):

h= height r=radius \[\LARGE d=2r\] d=diameter and you have height h=14 diameter d=9 can you find radius?

OpenStudy (anonymous):

how do you mean french...